Find the vertex of the parabola given by the function f(x) = 2x^2 + 3, expressing non-integer values in fractional form. State whether the parabola opens up or down
Answer by CharlesG2(834)

Find the vertex of the parabola given by the function f(x) = 2x^2 + 3, expressing non-integer values in fractional form. State whether the parabola opens up or down

f(x) = 2x^2 + 3 this is of form f(x) = ax^2 + bx + c,
where a = 2, b = 0, c = 3,
a = 2 is positive, so the parabola does open upwards


the vertex form of the parabolic equation is f(x) = a(x - h)^2 + k,
where the vertex is the point (h,k)

in this case h = 0, and k = c = 3

the vertex is (0, 3)